OpenStudy (anonymous):

solve this compound inequality? 4x>-20 and 6x<48

OpenStudy (anonymous):

Just isolate the x so you divide the first one by 4 and the second one by 6 x>-5 and x<8 since -20/4 =-5 and 48/6=8

OpenStudy (anonymous):

im terrible at this stuff :/ dont get it

OpenStudy (anonymous):

You have to get the 4 away from the x, and the only way to do that is by dividing it from both sides 4x/4 =x and -20/4=5 so the final inequality would be x>-5(answer) You would do the same for the other equation. 6x/6=6 and 48/6=8 so the final inequality would be x<8(answer)